DOPPLER-FREE TWO-PHOTON EXCITATION SPECTROSCOPY OF BENZENE IN A MOLECULAR BEAM

\underline{YOSUKE SEMBA} and SHUNJI KASAHARA, Molecular Photoscience Research Center, Kobe University, Kobe 657-8501, Japan; MASAAKI BABA, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan.

The fluorescence quantum yield at the vibrational energy over 3000 cm^{−1} in the S_{1} ^{1}B_{2u} state of benzene is dramatically decreased. This phenomenon is called ”channel three” and has been studied extensively. We have observed the ultrahighresolution spectrum of the S_{1} ← S_{0} 1_{0}^{2}14_{0}^{1} band ( E_{excess} = 3412 cm^{−1}) and found the Zeeman splitting which explained to be originating from the magnetic moment of the S_{1} state induced by mixing with S_{2} state by J-L coupling. ^{a} ^{b} In this work, we established the apparatus for the Doppler-free two-photon excitation (DFTPE) spectroscopy with a molecular beam and measured the S_{1} ← S_{0} 1_{0}^{1}14_{0}^{1} and 1_{0}^{2}14_{0}^{1} band of benzene. When we measured the DFTPE spectrum in a gas cell, many rotational lines were observed but the pressure broadening was remained. On the other hand, when we measured the DFTPE spectrum in a molecular beam, the influence of collision was decreased. The typical linewidth was 6 MHz. The high-resolution and collision-free condition is useful to investigate the perturbing region with high sensitivity and accuracy.

aD. Y. Baek, J. Chen, J. Wang, A. Doi, S. Kasahara, M. Baba, and H. Katô, Bul. Chem. Soc. Jpn. ,79, 75 (2006).

bH. Katô, M. Baba, and S. Kasahara, Bul. Chem. Soc. Jpn., 80, 456 (2007).